Dear sir, pLease help me on this problem, I got stuck,
1) A peacock is sitting on the top of a pillar which is 9m high .from a point 27m away from the bottom of pillar ,a snake is coming to its hole at the base of the pillar. seeing the snake the peacock pounces on it. if their speeds are equal ,at what distance from the hole is the snake caught?

Answer by scott8148(6628)

where x is the distance from the hole

the peacock travels ___ sqrt(9^2 + x^2)

the snake travels ___ 27 - x

sqrt(9^2 + x^2) = 27 - x

x^2 + 9^2 = x^2 - 54x + 27^2

54x = 27^2 - 9^2